# On Softmax Direct Preference Optimization for Recommendation

### v1

@@ Abstract
Recommender systems aim to predict personalized rankings based on user preference data. With the rise of Language Models (LMs), LM-based recommenders have been widely explored due to their extensive world knowledge and powerful reasoning abilities. Most of the LM-based recommenders convert historical interactions into language prompts, pairing with a positive item as the target response and fine-tuning LM with a language modeling loss. However, the current objective fails to fully leverage preference data and is not optimized for personalized ranking tasks, which hinders the performance of LM-based recommenders. Inspired by the current advancement of Direct Preference Optimization (DPO) in human preference alignment and the success of softmax loss in recommendations, we propose Softmax-DPO (S-DPO) to instill ranking information into the LM to help LM-based recommenders distinguish preferred items from negatives, rather than solely focusing on positives. Specifically, we incorporate multiple negatives in user preference data and devise an alternative version of DPO loss tailored for LM-based recommenders, connected to softmax sampling strategies. Theoretically, we bridge S-DPO with the softmax loss over negative sampling and find that it has a side effect of mining hard negatives, which assures its exceptional capabilities in recommendation tasks. Empirically, extensive experiments conducted on three real-world datasets demonstrate the superiority of S-DPO to effectively model user preference and further boost recommendation performance while mitigating the data likelihood decline issue of DPO. Our codes are available at https://github.com/chenyuxin1999/S-DPO.
@@ 1 Introduction
Recommender systems aim to predict personalized rankings based on user preference data, i.e., historical interactions such as purchases, clicks, and ratings [1, 2]. Recently, leveraging the extensive world knowledge and powerful reasoning abilities of language models (LMs) [3, 4, 5], LM-based recommenders have been broadly explored [6, 7, 8]. These recommenders convert historical interaction data into language prompts and either perform in-context learning or fine-tune LMs, demonstrating notable advantages, including zero-shot and few-shot reasoning [9, 10, 11, 12], enhanced generalization abilities [13, 14], and rich semantic understanding [15, 16, 17, 18]. However, current LM-based recommenders typically utilize language modeling loss for personalized ranking objectives—predicting the next token—which significantly differs from the objective of modeling user preferences in recommendation tasks [19, 20].
@@
Inspired by the success of using human-labeled data to align LMs with human preferences [30, 31, 32] and advancements in direct preference optimization (DPO) [33, 34, 35], we make progress on aligning LMs with recommendations by fine-tuning them to predict the next item in accordance with the user’s preference. This preference alignment stage aims to instill ranking information into the LMs and help recommenders distinguish preferred items from negatives, rather than solely focusing on positives.
Towards this end, we incorporate multiple negatives in user preference data and devise an alternative version of DPO loss tailored for recommendation, connected to softmax sampling strategies [19, 36, 37, 38], which we call S-DPO. Specifically, we first devise supervised fine-tuning to inject domain knowledge and improve LM’s ability to follow the instructions before preference alignment phase, following [13, 32]. In the preference alignment stage, instead of constructing solely positive pairs, we initially pair each language prompt with both positive and randomly sampled multiple negatives to build text-based preference data. Building upon these preference data, we extend conventional DPO with the Bradley-Terry preference model [33, 39] on pairwise data to the Plackett-Luce preference model [40, 41], which handles relative rankings in recommendation tasks.
Benefiting from the use of multiple negatives in preference data, our S-DPO offers three appealing properties. On the one hand, S-DPO serves as the first specialized personalized ranking loss for LM-based recommenders, effectively utilizing multiple negatives and acknowledging the importance of preference data. Empirically, we demonstrate that it provides more effective ranking gradients and mitigates the instability associated with DPO training (cf. Section 4.2). On the other hand, we theoretically bridge the DPO loss with the traditional BPR loss [42] over pairwise data and connect S-DPO with the softmax loss over negative sampling (also known as contrastive loss in self-supervised recommendations, which achieves state-of-the-art performance [36, 43, 44]). This connection naturally underscores the ranking performance of S-DPO and highlights the critical role of multiple negatives. Furthermore, we find that S-DPO has a side effect of mining hard negative examples similar to contrastive learning paradigm [37], which not only boosts the performance but also accelerates the training process (cf. Section 3.1), assuring its exceptional capabilities in recommendation tasks.
@@
  • We propose S-DPO, an alternative version of DPO loss tailored for LM-based recommenders, incorporating multiple negatives to instill ranking information into LM.
  • We theoretically bridge S-DPO with the softmax loss over negative sampling to highlight the critical role of multiple negatives and find its side effect of mining hard negatives, assuring its capabilities.
@@ 2 Preliminary
$\forall i_{d}\in\mathcal{I}_{d},\quad i_{p}>_{u}i_{d},$ (1)
wherein $>_{u}$ stands for the preference of user $u$.
@@
However, recommendation tasks are essentially user preference alignment tasks, as formalized in the above task formulation, and differ from user modeling tasks that consider only positive responses. Such a gap necessitates further exploration into aligning LM-based recommenders with user preference, an area that has been underexplored.
@@ 3.1 Derivation of S-DPO loss
As mentioned in Section 2, for user $u$, there is a partial ranking stipulating $i_{p}>_{u}i_{d},\forall\,i_{d}\in\mathcal{I}_{d}$ in sequential recommendation tasks. Let $\mathcal{E}_{d}$ be the titles of dispreferred items $\mathcal{I}_{d}$. The aforementioned partial ranking is equivalent to $e_{p}\succ e_{d}|x_{u},\forall e_{d}\in\mathcal{E}_{d}$, from which a multi-negative preference dataset $\{x_{u},e_{p},\mathcal{E}_{d}\}$ can be curated in an analogous way to RLHF.
For the dataset pairing one preferred item with multiple dispreferred items, we leverage the Plackett-Luce (PL) model [40, 41] to build preference distribution. Given prompt $x_{u}$, $K$ titles $e_{1},e_{2},\cdots,e_{K}$ and a permutation $\tau:[K]\rightarrow[K]$ reflecting user preference, with $\tau(j)$ denoting the $j$-th element of permutation $\tau$, the PL model estimates that the ranking $e_{\tau(1)},e_{\tau(2)},\cdots,e_{\tau(K)}$ turns out true, as:
@@
By substituting reward function $r(x_{u},e)$ in Eq.(9) with Eq.(6), the mutli-negative preference distribution can be rewritten as:
@@
Moreover, S-DPO contains an extra weight term only assigned to the gradient of dispreferred items. The term $\frac{1}{\sum_{e^{\prime}_{d}\in\mathcal{E}_{d}}{\rm exp}(g(e^{\prime}_{d},e_{d},x_{u}))}=\frac{{\rm exp}(r_{\theta}(x_{u},e_{d}))}{\sum_{e^{\prime}_{d}\in\mathcal{E}_{d}}{\rm exp}(r_{\theta}(x_{u},e^{\prime}_{d}))}$ mirrors the relative reward of item $i_{d}$ in dispreferred items. The larger the reward of item $i_{d}$ is compared with other dispreferred items, the higher the weight will be. Therefore, for dispreferred items with larger reward, which can be considered as hard negative items, their likelihood will decline more in the next update. This mechanism endows S-DPO with more effectiveness and stability than DPO. The gradient of S-DPO also tends to be larger, enabling S-DPO to converge faster .
@@ 3.2 Properties of S-DPO
In this section, we will discuss the structural correlation between DPO and BPR [42], together with S-DPO and softmax loss [38], which demonstrates the advantage of S-DPO over DPO and language modeling objective.
@@
Review the DPO loss in Eq.(7) and S-DPO loss in Eq.(11). Notably, term $\beta\,{\rm log}\,\frac{\pi_{\theta}(e|x_{u})}{\pi_{\rm ref}(e|x_{u})}$ is the implicit reward function, denoted by $r_{\theta}(x_{u},e)$ in Section 3.1. According to Section 2, $r_{\theta}(e,x_{u})$ reflects the preference of user $u$ to item $i$ corresponding to title $e$. When the reference model has no knowledge about recommendation, i.e., $\pi_{\rm ref}(e|x_{u})$ is approximately a uniform distribution, term $r_{\theta}(x_{u},e)=\beta\,{\rm log}\,\frac{\pi_{\theta}(e|x_{u})}{\pi_{\rm ref}(e|x_{u})}$ exactly reveals absolute preference. Hence, $r_{\theta}(x_{u},e)$ possesses a similar function to $f(u,i)$
From this perspective, DPO and S-DPO can be seen as special patterns of BPR and softmax loss, respectively. The structural correlation indicates that DPO and S-DPO are more suitable for recommendation than the language-modeling loss. Moreover, as softmax loss works better than BPR loss in multi-negative scenarios [19], it can be inferred that S-DPO will be more tailored for multi-negative user preference alignment than DPO.
@@ 4 Experiments
  • RQ2: How does the LM-based recommender benefit from the multiple negatives?
@@
We implement all LM-based recommenders on 4 NVIDIA A100 GPUs. For all LM-based recommenders, we conduct a supervised fine-tuning stage for a maximum of 5 epochs. For S-DPO and its variants, we conduct a preference alignment stage for further 3 epochs. Different from existing methods, we only optimize loss on item titles and find it effective in recommendation tasks. Refer to Appendix B for more implementation details.
@@ 4.1 Overall Performance Comparison (RQ1)
Table 1 presents a comparative analysis of the performance of our proposed S-DPO and baselines. “Rel.Ipv” denotes the relative improvement of S-DPO compared with baselines. Bold and underlined indicate the best and the second-best performance, respectively. We observe that:
@@
  • LM-based recommenders have driven impressive performance breakthroughs compared with traditional recommenders. Our results reveal that traditional recommenders outperform untuned LM-based recommenders (LLaMA, ChatRec) but fall short compared to LM-based recommenders fine-tuned on historical interactions (TALLRec and LLaRA). It is noted that untuned LM-based recommenders are limited by inadequate instruction-following capabilities or a lack of domain-specific knowledge indicated by the low valid ratio and suboptimal performance, which highlights the necessity of the supervised fine-tuning stage to further ground the inherent ability of language models down to sequential recommendation tasks. Moreover, MoRec also exhibits suboptimal performance compared to its traditional variant because it leaves the reasoning ability of LM untouched. The superior performance of recent LM-based recommenders indicates the significant roles of knowledge and reasoning ability in language models for recommendation tasks, which highlights the potential of LM-based recommenders.
@@
  • Tailoring language model for recommendation task further boosts the performance of LM-based recommenders. For LM-based recommenders, the substantial performance gap between fine-tuned and untuned approaches emphasizes the importance of tailoring models for recommendations. TALLRec adapts LM for recommendation by supervised fine-tuning LM on historical interactions, surpassing traditional recommenders. Additionally, LLaRA consistently outperformed TALLRec across all datasets, suggesting that introducing collaborative signals through appropriate item representations is a viable direction for further adapting LM. However, existing LM-based methods adapt LM from either item representation methods or corpus construction, leaving the adaptation of optimization objectives unexplored. Instead, S-DPO aligns the language model with multi-negative user preference data by extending DPO to include a softmax ranking loss, making it a more appropriate loss function for recommendation tasks.
@@
  • S-DPO consistently outperforms all traditional recommenders and the state-of-the-art LM-based recommenders on all datasets. S-DPO shows an improvement ranging from 11.10% to 47.03% on Hit Ratio@1 compared to the second-best baseline. Building on a supervised fine-tuning stage, we attribute this further improvement to the preference alignment stage, which explicitly instills ranking information into LM and utilizes preference data with multiple negative samples. Such superior performance suggests that explicitly tailoring LM for recommendation using user preference data at the training objective level is more effective than other LM-based recommenders. By leveraging the inherent abilities of the LM and incorporating ranking information from user preference data, S-DPO effectively differentiates between preferred and less preferred items. Notably, the preference alignment stage hardly harms the inherent ability of LM, illustrated by a high valid ratio.
Table 1: The performance comparison on three real-world datasets. The improvement achieved by S-DPO is significant ($p$-value <<much-less-than$<$ < 0.05).
LastFM Goodreads MovieLens
Traditional GRU4Rec [45] 0.3867 1.0000 70.91% 0.2616 1.0000 153.36% 0.3750 1.0000 40.35%
Caser [46] 0.4174 1.0000 58.34% 0.2233 1.0000 196.82% 0.3861 1.0000 36.31%
SASRec [47] 0.3581 1.0000 84.56% 0.2233 1.0000 196.82% 0.3444 1.0000 52.82%
LM-based LLaMA2 [31] 0.0233 0.3845 2736.48% 0.0246 0.3443 2594.31% 0.0421 0.4421 1150.12%
ChatRec [51] 0.3306 1.0000 99.91% 0.3770 1.0000 75.81% 0.2000 0.9895 163.15%
MoRec [48] 0.2877 1.0000 129.72% 0.1652 1.0000 301.21% 0.2822 1.0000 86.50%
TALLRec [13] 0.4983 0.9573 32.63% 0.4180 0.9836 58.56% 0.3895 0.9263 35.12%
LLaRA [18] 0.5292 0.9950 24.89% 0.4508 0.9918 47.03% 0.4737 0.9684 11.10%
Ours S-DPO 0.6609 0.9900 - 0.6628 0.9992 - 0.5263 0.9895 -
4.2 Study on S-DPO (RQ2)

### v3

@@ Abstract
Recommender systems aim to predict personalized rankings based on user preference data. With the rise of Language Models (LMs), LM-based recommenders have been widely explored due to their extensive world knowledge and powerful reasoning abilities. Most of the LM-based recommenders convert historical interactions into language prompts, pairing with a positive item as the target response and fine-tuning LM with a language modeling loss. However, the current objective fails to fully leverage preference data and is not optimized for personalized ranking tasks, which hinders the performance of LM-based recommenders. Inspired by the current advancement of Direct Preference Optimization (DPO) in human preference alignment and the success of softmax loss in recommendations, we propose Softmax-DPO (S-DPO) to instill ranking information into the LM to help LM-based recommenders distinguish preferred items from negatives, rather than solely focusing on positives. Specifically, we incorporate multiple negatives in user preference data and devise an alternative version of DPO loss tailored for LM-based recommenders, which is extended from the traditional full-ranking Plackett-Luce (PL) model to partial rankings and connected to softmax sampling strategies. Theoretically, we bridge S-DPO with the softmax loss over negative sampling and find that it has an inherent benefit of mining hard negatives, which assures its exceptional capabilities in recommendation tasks. Empirically, extensive experiments conducted on three real-world datasets demonstrate the superiority of S-DPO to effectively model user preference and further boost recommendation performance while providing better rewards for preferred items. Our codes are available at https://github.com/chenyuxin1999/S-DPO.
@@ 1 Introduction
Recommender systems aim to predict personalized rankings based on user preference data, i.e., historical interactions such as purchases, clicks, and ratings [1, 2]. Recently, leveraging the extensive world knowledge and powerful reasoning abilities of language models (LMs) [3, 4, 5, 6], LM-based recommenders have been broadly explored [7, 8, 9]. These recommenders convert historical interaction data into language prompts and either perform in-context learning or fine-tune LMs, demonstrating notable advantages, including zero-shot and few-shot reasoning [10, 11, 12, 13], enhanced generalization abilities [14, 15], and rich semantic understanding [16, 17, 18, 19]. However, current LM-based recommenders typically utilize language modeling loss for personalized ranking objectives—predicting the next token—which significantly differs from the objective of modeling user preferences in recommendation tasks [20, 21].
@@
Inspired by the success of using human-labeled data to align LMs with human preferences [31, 32, 33] and advancements in direct preference optimization (DPO) [34, 35, 36], we make progress on aligning LMs with recommendations by fine-tuning them to predict the next item in accordance with the user’s preference. This preference alignment stage aims to instill ranking information into the LMs and help recommenders distinguish preferred items from negatives, rather than solely focus on positives.
Towards this end, we incorporate multiple negatives in user preference data and devise an alternative version of DPO loss tailored for recommendation, connected to softmax sampling strategies [20, 37, 38, 39], which we call S-DPO. Specifically, we first devise supervised fine-tuning to inject domain knowledge and improve LM’s ability to follow the instructions before preference alignment phase, following [14, 33]. In the preference alignment stage, instead of constructing solely positive pairs, we initially pair each language prompt with both positive and randomly sampled multiple negatives to build text-based preference data. Building upon these preference data, we extend conventional DPO with the Bradley-Terry preference model [34, 40] on pairwise data to the Plackett-Luce preference model [41, 42], which handles relative rankings among multiple samples. Furthermore, we generalize the traditional Plackett-Luce preference model, which is designed for full relative rankings, to accommodate partial rankings, a more natural fit for recommendation tasks.
Benefiting from the multiple negatives in preference data, S-DPO offers three appealing properties. On the one hand, S-DPO serves as the first specialized personalized ranking loss for LM-based recommenders, effectively utilizing multiple negatives and acknowledging the importance of preference data. Empirically, we demonstrate that it provides more effective ranking gradients and better rewards for preferred items compared with DPO (cf. Section 4.2). On the other hand, we theoretically bridge the DPO loss with the pairwise BPR loss [43] over pairwise data and connect S-DPO with the softmax loss over negative sampling (also known as contrastive loss in self-supervised recommendations, which achieves state-of-the-art performance [37, 44, 45]). This connection naturally underscores the ranking performance of S-DPO and highlights the critical role of multiple negatives. Furthermore, gradient analysis demonstrates that S-DPO has an inherent benefit of mining hard negative examples similar to contrastive learning paradigm [38], which not only boosts the performance but also accelerates the training process (cf. Section 3.1), assuring its exceptional capabilities in recommendation tasks.
@@
  • We propose S-DPO, an alternative version of DPO loss extended from the traditional Plackett-Luce preference model, incorporating multiple negatives to instill ranking information into LM and tailoring for LM-based recommenders.
  • We theoretically bridge S-DPO with the softmax loss over negative sampling to highlight the critical role of multiple negatives and find its inherent benefit of mining hard negatives, assuring its capabilities.
@@ 2 Preliminary
$\forall i_{d}\in\mathcal{I}_{d},\quad i_{p}\succ_{u}i_{d},$ (1)
wherein $\succ_{u}$ stands for the preference of user $u$.
@@
However, recommendation tasks are essentially user preference alignment tasks, as formalized in the above task formulation, and differ from language modeling tasks that consider only positive responses. Such a gap necessitates further exploration into aligning LM-based recommenders with user preference, an area that has been underexplored.
@@ 3.1 Derivation of S-DPO loss
As mentioned in Section 2, for user $u$, there is a partial ranking stipulating $i_{p}\succ_{u}i_{d},\forall\,i_{d}\in\mathcal{I}_{d}$ in sequential recommendation tasks. Let $\mathcal{E}_{d}$ be the titles of dispreferred items $\mathcal{I}_{d}$. The aforementioned partial ranking is equivalent to $e_{p}\succ e_{d}|x_{u},\forall e_{d}\in\mathcal{E}_{d}$, from which a multi-negative preference dataset $\{x_{u},e_{p},\mathcal{E}_{d}\}$ can be curated in an analogous way to RLHF.
For the dataset pairing one preferred item with multiple dispreferred items, we leverage the Plackett-Luce (PL) model [41, 42] to build preference distribution. Given prompt $x_{u}$, $K$ titles $e_{1},e_{2},\cdots,e_{K}$ and a permutation $\tau:[K]\rightarrow[K]$ reflecting the user preference, with $\tau(j)$ denoting the $j$-th element of permutation $\tau$, the PL model estimates that the ranking $e_{\tau(1)},e_{\tau(2)},\cdots,e_{\tau(K)}$ turns out true, as:
@@
By substituting reward function $r(x_{u},e)$ in Eq.(9) with Eq.(6), the multi-negative preference distribution can be rewritten as:
@@
Moreover, S-DPO treats gradients of different negative (dispreferred) items differently by assigning the gradient of each negative item with an extra weight $\frac{1}{\sum_{\epsilon^{\prime}_{d}\in\mathcal{E}_{d}}{\rm exp}(g(e^{\prime}_{d},e_{d},x_{u}))}=\frac{{\rm exp}(r_{\theta}(x_{u},e_{d}))}{\sum_{\epsilon^{\prime}_{d}\in\mathcal{E}_{d}}{\rm exp}(r_{\theta}(x_{u},e^{\prime}_{d}))}$. This term reflects the relative reward of each negative item compared with other negative items. Similar to [38], we can categorize negative items into two categories: (1) Hard negative items, whose reward $r_{\theta}(x_{u},e_{d})=\beta\frac{\pi_{\theta}(e_{d}|x_{u})}{\pi_{\rm ref}(e_{d}|x_{u})}$ is relatively high, making it more probable to be chosen by LM-based recommenders; (2) Easy negative items, whose reward $r_{\theta}(x_{u},e_{d})$ is relatively low, making it less likely to be output. For hard negative items, the extra weight term $\frac{{\rm exp}(r_{\theta}(x_{u},e_{d}))}{\sum_{\epsilon^{\prime}_{d}\in\mathcal{E}_{d}}{\rm exp}(r_{\theta}(x_{u},e^{\prime}_{d}))}$ tends to be larger, leading to more decline for likelihood. This mechanism makes LM-based recommenders more discriminative and endows S-DPO with more effectiveness and stability than DPO.
@@ 3.2 Properties of S-DPO
In this section, we will discuss the structural correlation between DPO and BPR [43], together with S-DPO and softmax loss [39], which demonstrates the advantage of S-DPO over DPO and language modeling loss.
@@
Review the DPO loss in Eq.(7) and S-DPO loss in Eq.(11). Notably, term $\beta\,{\rm log}\,\frac{\pi_{\theta}(e|x_{u})}{\pi_{\rm ref}(e|x_{u})}$ is the implicit reward function, denoted by $r_{\theta}(x_{u},e)$ in Section 3.1. According to Section 2, $r_{\theta}(e,x_{u})$ reflects the preference of user $u$ to item $i$ corresponding to title $e$. When the reference model has no knowledge about recommendation, i.e., when $\pi_{\rm ref}(e|x_{u})$ is approximately a uniform distribution, term $r_{\theta}(x_{u},e)=\beta\,{\rm log}\,\frac{\pi_{\theta}(e|x_{u})}{\pi_{\rm ref}(e|x_{u})}$ exactly reveals absolute preference. Hence, $r_{\theta}(x_{u},e)$ possesses a similar function to $f(u,i)$
From this perspective, DPO and S-DPO can be seen as special patterns of BPR and softmax loss, respectively. Given the effectiveness of BPR and InfoNCE loss in recommendation, we argue that sampled-based loss which explicitly compares preferred and dispreferred items such as DPO and S-DPO is more suitable for training LM-based recommenders than only utilizing language modeling loss. Moreover, as softmax loss works better than BPR loss in multi-negative scenarios [20], it can be inferred that S-DPO will be more tailored for multi-negative user preference alignment than DPO.
@@ 4 Experiments
  • RQ2: How does the LM-based recommender benefit from the multiple negatives in S-DPO?
  • RQ3: What are the impacts of the essential parameters ($\beta$) on S-DPO?
@@
Table 1: The performance comparison on three real-world datasets. “Rel.Ipv” denotes the relative improvement of S-DPO compared with baselines.
Goodreads LastFM MovieLens
Traditional GRU4Rec 0.3867 1.0000 70.91% 0.2616 1.0000 153.36% 0.3750 1.0000 40.35%
Caser 0.4174 1.0000 58.34% 0.2233 1.0000 196.82% 0.3861 1.0000 36.31%
SASRec 0.3581 1.0000 84.56% 0.2233 1.0000 196.82% 0.3444 1.0000 52.82%
LM-based LLaMA2 0.0233 0.3845 2736.48% 0.0246 0.3443 2594.31% 0.0421 0.4421 1150.12%
ChatRec 0.3306 1.0000 99.91% 0.3770 1.0000 75.81% 0.2000 0.9895 163.15%
MoRec 0.2877 1.0000 129.72% 0.1652 1.0000 301.21% 0.2822 1.0000 86.50%
TALLRec 0.4983 0.9573 32.63% 0.4180 0.9836 58.56% 0.3895 0.9263 35.12%
LLaRA 0.5292 0.9950 24.89% 0.4508 0.9918 47.03% 0.4737 0.9684 11.10%
Ours S-DPO 0.6609 0.9900 - 0.6628 0.9992 - 0.5263 0.9895 -
@@
We implement all LM-based recommenders on 4 NVIDIA A100 GPUs. For all LM-based recommenders, we conduct a supervised fine-tuning stage for a maximum of 5 epochs. For S-DPO and its variants, we conduct a preference alignment stage for a further 3 epochs. Different from existing methods, we only optimize loss on item titles and find it effective in recommendation tasks. Refer to Appendix B for more implementation details.
@@ 4.1 Overall Performance Comparison (RQ1)
Table 1 presents a comparative analysis of the performance of our proposed S-DPO and baselines. Bold and underlined indicate the best and the second-best performance, respectively. We observe that:
@@
  • LM-based recommenders have driven impressive performance breakthroughs compared with traditional recommenders. Our results reveal that traditional recommenders outperform untuned LM-based recommenders (LLaMA, ChatRec) but fall short compared to LM-based recommenders fine-tuned on historical interactions (TALLRec and LLaRA). It is noted that untuned LM-based recommenders are limited by inadequate instruction-following capabilities (indicated by a low valid ratio) or a lack of domain-specific knowledge (indicated by a suboptimal performance), which highlights the necessity of the supervised fine-tuning stage to further ground the inherent ability of language models down to sequential recommendation tasks. Moreover, MoRec also exhibits suboptimal performance compared to its traditional variant because it leaves the reasoning ability of LM untouched. The superior performance of recent LM-based recommenders indicates the significant roles of knowledge and reasoning ability in language models for recommendation tasks in semantically informative datasets, which highlights the potential of LM-based recommenders.
@@
  • Tailoring language models for recommendation task further boosts the performance of LM-based recommenders. For LM-based recommenders, the substantial performance gap between fine-tuned and untuned approaches emphasizes the importance of tailoring models for recommendations. TALLRec adapts LM for recommendation by supervised fine-tuning LM on historical interactions, surpassing traditional recommenders. Additionally, LLaRA consistently outperformed TALLRec across all datasets, suggesting that introducing collaborative signals through appropriate item representations is a viable direction for further adapting LM. However, existing LM-based methods adapt LM from either item representation methods or corpus construction, leaving the adaptation of optimization objectives unexplored. Instead, S-DPO aligns the language model with multi-negative user preference data by extending DPO to include a softmax ranking loss, making it a more appropriate loss function for recommendation tasks.
@@
  • S-DPO consistently outperforms all traditional recommenders and the state-of-the-art LM-based recommenders on all datasets. S-DPO shows an improvement ranging from 11.10% to 47.03% on Hit Ratio@1 compared to the second-best baseline. Building on a supervised fine-tuning stage, we observe a further improvement to the preference alignment stage, which explicitly instills ranking information into LM and utilizes preference data with multiple negative samples. Such superior performance suggests that explicitly tailoring LM for recommendation using user preference data at the training objective level is more effective than other LM-based recommenders. By leveraging the inherent abilities of the LM and incorporating ranking information from user preference data, S-DPO effectively differentiates between preferred and less preferred items. Notably, the preference alignment stage hardly harms the inherent ability of LM, illustrated by a high valid ratio.
4.2 Study on S-DPO